find the measures of the angles of a triangle if the measure of one side is twice the measure of a second and the third angle measures two times the second decreased by 20
Answer by lwsshak3(11628)

find the measures of the angles of a triangle if the measure of one side is twice the measure of a second and the third angle measures two times the second decreased by 20
***
let x=2nd angle
2x=1st angle
2x-20=3rd angle
..
x+2x+2x-20=180
5x=200
x=40
2x=80
2x-20=60
..
measure of 2nd angle=40˚
measure of 1st angle=80˚
measure of 3rd angle=60˚